# Measurement of $\Sigma^{+}$ transverse polarization in $e^{+}e^{-}$ collisions at $\sqrt{s}=3.68-3.71$ GeV

### v2

@@ Abstract
Using $e^{+}e^{-}$ collision data collected with the BESIII detector at seven energy points ranging from 3.68 to 3.71 GeV and corresponding to an integrated luminosity of $652.1~{}{\rm pb^{-1}}$, we present an energy-dependent measurement of the transverse polarization, relative phase and modulus ratio of the electromagnetic form factors of the $\Sigma^{+}$ hyperon in the $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ reaction. These results are helpful to understand the production mechanism of the $\Sigma^{+}$-$\bar{\Sigma}^{-}$ pairs.
@@ 1 Introduction
In accordance with the Phragmén-Lindelöf theorem titchmarsh1939theory , as $q^{2}\to\infty$, the asymptotic behavior of the time-like EMFFs can be obtained from their corresponding space-like counterparts. In the space-like region, the EMFFs are assumed to be real. Consequently, as $q^{2}\to\infty$, the time-like EMFFs must also be real, and the relative phase of the EMFFs should approach an integer multiple of $\pi$. In the case where $q^{2}$ does not approach the limit, the phase can be any value other than zero, which causes the final state to be polarized, even if the initial state is unpolarized Dubnickova:1992ii . This allows the determination of the EMFFs, as the polarization and cross-section are functions of form factors themselves Faldt:2017kgy ; Cao:2021asd ; Dai:2023vsw . Recent results from BESIII BESIII:2020uqk , BaBar Aubert:2007uf and Belle Belle:2022dvb , focusing solely on measuring the effective form factor, have reported consistent results. Additionally, the CLEO collaboration has made significant progress in determining the cross sections and the EMFFs of various baryon pairs ($p$, $\Lambda$, $\Sigma^{0}$, $\Sigma^{+}$, $\Xi^{0}$, $\Xi^{-}$, and $\Omega^{-}$) Dobbs:2014ifa ; Dobbs:2017 . Their conclusions regarding EMFFs and di-quark correlations Jafee:2005 ; Jafee:2003 rely on the assumption that one-photon exchange dominates the production process and that charmonia contributions are negligible. Furthermore, the BESIII collaboration has measured the cross sections for several baryon pairs ($\Lambda$, $\Sigma^{0}$, $\Xi^{-}$, $\Sigma^{\pm}$, and $\Xi^{0}$) near the production threshold BESIII:2017hyw ; BESIII:2020uqk ; BESIII:2020ktn ; BESIII:2021aer ; BESIII:2021rkn ; BESIII:2022kzc and above open charm threshold BESIII:2019cuv ; BESIII:2021ccp ; BESIII:2023rse ; BESIII:2024umc ; BESIII:2024ogz . However, experimental measurements of the relative phase between $G_{E}$ and $G_{M}$ are still relatively scarce.
@@
For spin 1/2 baryon pairs produced by the $e^{+}e^{-}$ process via vector charmonia, where the above statements concerning EMFFs also apply, the formalism is described in ref. Faldt:2017kgy . In these cases, the amplitudes include the EM-psionic form factors, $G^{\Psi}_{E}$ and $G^{\Psi}_{M}$ BESIII:2021cvv ; BESIII:2023euh . While the EM-psionic form factors describe a different process, the form of the hadron current matrix element for the charmonia process is the same as that for the virtual photon one Faldt:2017kgy ; Faldt:EPJ . Previous studies often neglected the polarization effects of hyperons Ablikim:2016iym ; Ablikim:2016iym-01 ; Ablikim:2016iym-02 ; Ablikim:2016iym-03 ; Wang:2018kdh ; Wang:2021lfq . However, $\Sigma^{+}$ polarization was recently observed and measured in the $e^{+}e^{-}\to J/\psi$, $\psi(3686)\to\Sigma^{+}\bar{\Sigma}^{-}$ processes by the BESIII collaboration BESIII:2020fqg ; BESIII:2023sgt . The results not only reveal a non-zero relative phase but also demonstrate that the phase changes sign at the mass of the $\psi(3686)$ resonance compared to the value measured at the $J/\psi$ resonance. Subsequently, the polarization effects of $\Lambda$ and $\Xi^{-}$ were also observed and measured by the BESIII collaboration. The $\Lambda$ polarization was observed in the $e^{+}e^{-}\to\Lambda\bar{\Lambda}$ process at the $J/\psi$, $\psi(3770)$, and off-resonance regions BESIII:2021cvv ; BESIII:2023euh ; BESIII:2022yprl ; Ablikim:2018zay . Additionally, a non-zero polarization has been observed for $\Xi^{-}$ in the $e^{+}e^{-}\rightarrow J/\psi$, $\psi(3686)\rightarrow\Xi^{-}\bar{\Xi}^{+}$ processes BESIII:2021ypr ; BESIII:2022_xi_psip ; Liu:2023xhg . However, for $\Xi^{0}$, completely different effects were observed in processes $e^{+}e^{-}\rightarrow J/\psi\rightarrow\Xi^{0}\bar{\Xi}^{0}$ BESIII:2023drj and $e^{+}e^{-}\rightarrow\psi(3686)\rightarrow\Xi^{0}\bar{\Xi}^{0}$ BESIII:2023lkg . The data samples used in this analysis correspond to an integrated luminosity of 652.1 pb-1, collected at the center-of-mass (CM) energies of $\sqrt{s}=$ 3.682, 3.683, 3.684, 3.685, 3.687, 3.691, and 3.710 GeV with the BESIII detector Ablikim:2009aa in symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring BEPCII . The above energy points around $\psi(3686)$ resonance, which are used to study the energy dependence of $\Delta\Phi$, are particularly intriguing, as the production occurs through a combination of one-photon exchange BESIII:2019nep , mixing with the $\psi(3686)$ resonance BESIII:2021cvv , and resonance dominance Ablikim:2018zay ; BESIII:2022yprl . Since the energy range studied here includes both the virtual photon vector charmonium processes, $G_{E}^{\gamma/\Psi}$ and $G_{M}^{\gamma/\Psi}$ are used to represent the form factors in the following, and the methods used in the following apply to both processes.
@@ 3 Event selection
Charged tracks detected in the MDC are required to be within a polar angle ($\theta$) range of $|\!\cos\theta|<0.93$, where $\theta$ is defined with respect to the $z$-axis, which is the symmetry axis of the MDC. At least two oppositely charged tracks, which must be well reconstructed in the MDC with good helix fits, are required. There is no vertex requirement. Since (anti-)protons from $\Sigma^{+}(\bar{\Sigma}^{-})$ decays can be distinguished from other charged particles by requiring the momentum to be greater than $0.5~{}{\rm GeV}/c$, there is no additional particle identification (PID) requirement.
@@ 4 $\Sigma^{+}$ polarization
The exclusive process $e^{+}e^{-}\rightarrow\gamma^{*}/\Psi\rightarrow\Sigma^{+}\bar{\Sigma}^{-}\rightarrow p\bar{p}\pi^{0}\pi^{0}$ can be fully described by the $\Sigma^{+}$ scattering angle, $\theta_{\Sigma^{+}}$, in the CM system of the $e^{+}e^{-}$ reaction and the $p$ and  $\bar{p}$ directions, $\bm{\hat{n}_{1}}$ and $\bm{\hat{n}_{2}}$, respectively, in the rest frames of their parent particles. Here $\gamma^{*}/\Psi$ indicates that the process $e^{+}e^{-}\rightarrow\Sigma^{+}\bar{\Sigma}^{-}$ occurs via a pure EM process or a $\psi$ resonance. The components of these vectors are expressed using a coordinate system ($x_{\Sigma^{+}},y_{\Sigma^{+}},z_{\Sigma^{+}}$) as shown in figure 2. A right-handed system for each hyperon decay is defined with the $z$ axis along the $\Sigma^{+}$ momentum $_{\Sigma^{+}}=-\textbf{p}_{\bar{\Sigma}^{-}}=\textbf{p}$ p in the CM system. The $y$ axis is taken as the normal to the scattering plane, $\textbf{k}_{e^{-}}\times\textbf{p}_{\Sigma^{+}}$, where $_{e^{-}}=-\textbf{k}_{e^{+}}=\textbf{k}$ k is the electron beam momentum in the CM system.
@@
${\cal{W}}(\bm{\xi};\bm{\Omega})=$ $\mathcal{F}_{0}(\xi)+\eta\mathcal{F}_{5}(\xi)+\alpha_{\Sigma^{+}}\alpha_{\bar{\Sigma}^{-}}$
$\times$ $[\mathcal{F}_{1}(\xi)+\sqrt{1-\eta^{2}}\cos(\Delta\Phi)\mathcal{F}_{2}(\xi)+\eta\mathcal{F}_{6}(\xi)]$
$+$ $\sqrt{1-\eta^{2}}\sin(\Delta\Phi)[\alpha_{\Sigma^{+}}\mathcal{F}_{3}(\xi)+\alpha_{\bar{\Sigma}^{-}}\mathcal{F}_{4}(\xi)],$ (1)
[…]
@@
$\mathscr{S}=-\mathrm{ln}\mathscr{L}_{\rm data}$ (6)
@@
Figure 3 shows the distributions of the five ${F}_{k}(k=1,2,...,5)$ moments Faldt:2017kgy with respect to $\cos\theta_{\Sigma^{+}}$, divided into 10 intervals, and the weighted PHSP MC results corrected by the global fit. The numerical fit results and the weighted average values are summarized in table 1. Here, the weighted average is calculated by $\frac{1}{w}\sum_{i=1}^{n}w_{i}x_{i}$ PDG2020 , where $n$ is the number of energy points, $w_{i}=1/\sigma_{i}^{2}$, $x_{i}$ and $\sigma_{i}$ are the measured values and their uncertainties at each energy point, respectively, and $w=\sum_{i}w_{i}$. The relative phases of the form factors at each energy point are measured to be different from zero with significances of less than $3\sigma$. The significance is estimated by comparing the likelihoods of the baseline fit and the one defined assuming no polarization p:value . The calculation of significance also accounts for the systematic uncertainties in the decay parameter, mass window, and background. After evaluating the significance following each systematic uncertainty, we select the smallest value as a conservative estimation. The significance of each energy point is also summarized in table 1.
Figure 3: Distributions of ${F}_{k}(k=1,2,...,5)$ moments with respect to $\cos\theta_{\Sigma^{+}}$. The dots with error bars are data combined from all energy points, and the red solid lines are the weighted PHSP MC corrected by the results of global fit. The blue dashed lines represent the distributions of the simulated events evenly distributed in PHSP, without polarization.
@@
$M(\cos\theta_{\Sigma^{+}})=\frac{\alpha_{\Sigma^{+}}-\alpha_{\bar{\Sigma}^{-}}}{2}\frac{1+\eta\cos^{2}\theta_{\Sigma^{+}}}{3+\eta}P_{y}(\theta_{\Sigma^{+}})$ (8)

### v3

@@ Abstract
Using $e^{+}e^{-}$ collision data collected with the BESIII detector at seven energy points ranging from 3.68 to 3.71 GeV and corresponding to an integrated luminosity of $652.1~{}{\rm pb^{-1}}$, we present an energy-dependent measurement of the transverse polarization, relative phase and modulus ratio of the electromagnetic form factors of the $\Sigma^{+}$ hyperon in the $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ reaction. At each energy point, no evident polarization is observed. These results are helpful to understand the production mechanism of the $\Sigma^{+}$-$\bar{\Sigma}^{-}$ pairs.
@@ 1 Introduction
In accordance with the Phragmén-Lindelöf theorem titchmarsh1939theory , as $q^{2}\to\infty$, the asymptotic behavior of the time-like EMFFs can be obtained from their corresponding space-like counterparts. In the space-like region, the EMFFs are assumed to be real. Consequently, as $q^{2}\to\infty$, the time-like EMFFs must also be real, and the relative phase of the EMFFs should approach an integer multiple of $\pi$. In the case where $q^{2}$ does not approach the limit, the phase can be any value, which causes the final state to be polarized, even if the initial state is unpolarized Dubnickova:1992ii . This allows the determination of the EMFFs, as the polarization and cross-section are functions of form factors themselves Faldt:2017kgy ; Cao:2021asd ; Dai:2023vsw ; Yan:2023yff ; Chen:2024luh . Recent results from BESIII BESIII:2020uqk , BaBar Aubert:2007uf and Belle Belle:2022dvb , focusing solely on measuring the effective form factor, have reported consistent results. Additionally, the CLEO collaboration has made significant progress in determining the cross sections and the EMFFs of various baryon pairs ($p$, $\Lambda$, $\Sigma^{0}$, $\Sigma^{+}$, $\Xi^{0}$, $\Xi^{-}$, and $\Omega^{-}$) Dobbs:2014ifa ; Dobbs:2017 . Their conclusions regarding EMFFs and di-quark correlations Jafee:2005 ; Jafee:2003 rely on the assumption that one-photon exchange dominates the production process and that charmonia contributions are negligible. Furthermore, the BESIII collaboration has measured the cross sections for several baryon pairs ($\Lambda$, $\Sigma^{0}$, $\Xi^{-}$, $\Sigma^{\pm}$, and $\Xi^{0}$) near the production threshold BESIII:2017hyw ; BESIII:2020uqk ; BESIII:2020ktn ; BESIII:2021aer ; BESIII:2021rkn ; BESIII:2022kzc and above open charm threshold BESIII:2019cuv ; BESIII:2021ccp ; BESIII:2023rse ; BESIII:2024umc ; BESIII:2024ogz . However, experimental measurements of the relative phase between $G_{E}$ and $G_{M}$ are still relatively scarce.
For spin 1/2 baryon pairs produced by the $e^{+}e^{-}$ process via vector charmonia, where the above statements concerning EMFFs also apply, the formalism is described in ref. Faldt:2017kgy . In these cases, the amplitudes include the EM-psionic form factors, $G^{\Psi}_{E}$ and $G^{\Psi}_{M}$ BESIII:2021cvv ; BESIII:2023euh . While the EM-psionic form factors describe a different process, the form of the hadron current matrix element for the charmonia process is the same as that for the virtual photon one Faldt:2017kgy ; Faldt:EPJ .
Previous studies often neglected the polarization effects of hyperons Ablikim:2016iym ; Ablikim:2016iym-01 ; Ablikim:2016iym-02 ; Ablikim:2016iym-03 ; Wang:2018kdh ; Wang:2021lfq . However, $\Sigma^{+}$ polarization was recently observed and measured in the $e^{+}e^{-}\to J/\psi$, $\psi(3686)\to\Sigma^{+}\bar{\Sigma}^{-}$ processes by the BESIII collaboration BESIII:2020fqg ; BESIII:2023sgt . The results not only reveal a non-zero relative phase but also demonstrate that the phase changes sign at the mass of the $\psi(3686)$ resonance compared to the value measured at the $J/\psi$ resonance. Subsequently, the polarization effects of $\Lambda$ and $\Xi^{-}$ were also observed and measured by the BESIII collaboration. The $\Lambda$ polarization was observed in the $e^{+}e^{-}\to\Lambda\bar{\Lambda}$ process at the $J/\psi$, $\psi(3770)$, and off-resonance regions BESIII:2022yprl ; Ablikim:2018zay ; Wang:2023trb . Additionally, a non-zero polarization has been observed for $\Xi^{-}$ in the $e^{+}e^{-}\rightarrow J/\psi$, $\psi(3686)\rightarrow\Xi^{-}\bar{\Xi}^{+}$ processes BESIII:2021ypr ; BESIII:2022_xi_psip ; Liu:2023xhg . However, for $\Xi^{0}$, completely different effects were observed in processes $e^{+}e^{-}\rightarrow J/\psi\rightarrow\Xi^{0}\bar{\Xi}^{0}$ BESIII:2023drj and $e^{+}e^{-}\rightarrow\psi(3686)\rightarrow\Xi^{0}\bar{\Xi}^{0}$ BESIII:2023lkg .
@@
The data samples used in this analysis correspond to an integrated luminosity of 652.1 pb-1, collected at the center-of-mass (CM) energies of $\sqrt{s}=$ 3.682, 3.683, 3.684, 3.685, 3.687, 3.691, and 3.710 GeV with the BESIII detector Ablikim:2009aa in symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring BEPCII . The above energy points around $\psi(3686)$ resonance, which are used to study the energy dependence of $\Delta\Phi$, are particularly intriguing, as the production occurs through a combination of one-photon exchange BESIII:2019nep , mixed with the $\psi(3686)$ resonance BESIII:2021cvv , and resonance dominance Ablikim:2018zay ; BESIII:2022yprl . Also, it can enrich the experimental information of form factors measurement in higher energy region Ramalho:2024wxp . Since the energy range studied here includes both the virtual photon and vector charmonium processes, $G_{E}^{\gamma/\Psi}$ and $G_{M}^{\gamma/\Psi}$ are used to represent the form factors in the following, and the methods used in the following apply to both processes.
@@ 3 Event selection
Charged tracks detected in the MDC are required to be within a polar angle ($\theta$) range of $|\!\cos\theta|<0.93$, where $\theta$ is defined with respect to the $z$-axis, which is the symmetry axis of the MDC. At least two oppositely charged tracks, which must be well reconstructed in the MDC with good helix fits, are required. For the distance $\rm V_{xy}$ between the vertex and the xy plane, and the distance $\rm V_{z}$ between the vertex and the z-axis, both the signal and background distributions are around 0. To avoid efficiency loss, there is no vertex requirement. Since (anti-)protons from $\Sigma^{+}(\bar{\Sigma}^{-})$ decays can be distinguished from other charged particles by requiring the momentum to be greater than $0.5~{}{\rm GeV}/c$, there is no additional particle identification (PID) requirement.
@@ 4 $\Sigma^{+}$ polarization
The exclusive process $e^{+}e^{-}\rightarrow\gamma^{*}/\Psi\rightarrow\Sigma^{+}\bar{\Sigma}^{-}\rightarrow p\bar{p}\pi^{0}\pi^{0}$ can be fully described by the $\Sigma^{+}$ scattering angle, $\theta_{\Sigma^{+}}$, in the CM system of the $e^{+}e^{-}$ reaction and the $p$ and  $\bar{p}$ directions, $\bm{\hat{n}_{1}}$ and $\bm{\hat{n}_{2}}$, respectively, in the rest frames of their parent particles. Here $\gamma^{*}/\Psi$ indicates that the process $e^{+}e^{-}\rightarrow\Sigma^{+}\bar{\Sigma}^{-}$ occurs via a pure EM process or a $\psi$ resonance. The components of these vectors are expressed using a coordinate system ($x_{\Sigma^{+}},y_{\Sigma^{+}},z_{\Sigma^{+}}$) as shown in figure 2. A right-handed system for each hyperon decay is defined with the $z$ axis along the $\Sigma^{+}$ momentum $_{\Sigma^{+}}=-\textbf{p}_{\bar{\Sigma}^{-}}=\textbf{p}$ p in the CM system. The $y$ axis is taken as the normal to the scattering plane, $\textbf{p}_{\Sigma^{+}}\times\textbf{k}_{e^{-}}$, where $_{e^{-}}=-\textbf{k}_{e^{+}}=\textbf{k}$ k is the electron beam momentum in the CM system.
@@
${\cal{W}}(\bm{\xi};\bm{\Omega})=$ $F_{0}(\xi)+\eta F_{5}(\xi)+\alpha_{\Sigma^{+}}\alpha_{\bar{\Sigma}^{-}}$
$\times$ $[F_{1}(\xi)+\sqrt{1-\eta^{2}}\cos(\Delta\Phi)F_{2}(\xi)+\eta F_{6}(\xi)]$
$+$ $\sqrt{1-\eta^{2}}\sin(\Delta\Phi)[\alpha_{\Sigma^{+}}F_{3}(\xi)+\alpha_{\bar{\Sigma}^{-}}F_{4}(\xi)],$ (1)
where $\eta$ is the angular distribution parameter, $\Omega$ $=(\eta,\Delta\Phi,\alpha_{\Sigma^{+}},\alpha_{\bar{\Sigma}^{-}})$ represents the production and decay parameters, the kinematic variables $\xi$ $=(\theta_{\Sigma^{+}},\bm{\hat{n}_{1}},\bm{\hat{n}_{2}})$ describe the production and subsequent decay, and $\alpha_{\Sigma^{+}(\bar{\Sigma}^{-})}$ denotes the asymmetry parameter of the $\Sigma^{+}(\bar{\Sigma}^{-})\rightarrow p\pi^{0}(\bar{p}\pi^{0})$ decay. The angular functions $F_{i}(\xi)~{}(i=0,1,...6)$ are described in detail in ref. Faldt:2017kgy , and the definitions of five terms with angular distribution of final states are shown in eq. (4).
@@
$\mathscr{S}=-\mathrm{ln}\mathscr{L}_{\rm data},$ (6)
@@
Figure 3 shows the distributions of the five ${F}_{k}(k=1,2,3,4,6)$ moments Faldt:2017kgy with respect to $\cos\theta_{\Sigma^{+}}$, divided into 10 intervals, and the weighted PHSP MC results corrected by the global fit. The numerical fit results and the weighted average values are summarized in table 1. The weighted average can provide an overall evaluation of the $\Sigma^{+}$ polarization related to the production mechanism of the $\Sigma^{+}\bar{\Sigma}^{-}$ pair around this energy range. It is calculated by $\frac{1}{w}\sum_{i=1}^{n}w_{i}x_{i}$ PDG2020 , where $n$ is the number of energy points, $w_{i}=1/\sigma_{i}^{2}$, $x_{i}$ and $\sigma_{i}$ are the measured values and their uncertainties at each energy point, respectively, and $w=\sum_{i}w_{i}$. The significance is estimated by comparing the likelihoods of the baseline fit and the one defined assuming no polarization p:value . The calculation of significance also accounts for the systematic uncertainties in the decay parameter, mass window, and background. After evaluating the significance following each systematic uncertainty, we select the smallest value as a conservative estimation. The significance of each energy point is also summarized in table 1. Since the significances are all less than $3\sigma$, no evident polarization is observed. Note that the significance of polarization $P_{y}$ can not be directly determined by $\Delta\Phi$ and it should be directly determined by $\sin(\Delta\Phi)$ when $P_{y}\propto\sin(\Delta\Phi)=0$ according to eq. (4). It although means that $\Delta\Phi$ allows two values either $\Delta\Phi$ or $\Delta\Phi-\pi$, the terms $\sin(\Delta\Phi)$ and $\cos(\Delta\Phi)$ in eq.(4) can constraint the unique $\Delta\Phi$ value.
Figure 3: Distributions of ${F}_{k}(k=1,2,3,4,6)$ moments with respect to $\cos\theta_{\Sigma^{+}}$. The dots with error bars are data combined from all energy points, and the red solid lines are the weighted PHSP MC corrected by the results of global fit. The blue dashed lines represent the distributions of the simulated events evenly distributed in PHSP, without polarization.
[…]
$3682.00\pm 0.20$ 1.334 404.0 $134$. $2^{+12.4}_{-11.8}$ $0.54\pm 0.17\pm 0.12$ $0.38\pm 0.40\pm 0.12$ $0.84\pm 0.20\pm 0.14$ 0.9
$3682.75\pm 0.20$ 1.335 28.7 $27$. $2^{+6.0}_{-5.4}$ $0.96\pm 0.13\pm 0.12$ $2.35\pm 1.66\pm 0.12$ $0.22\pm 0.37\pm 0.34$ 0.5
$3684.22\pm 0.20$ 1.336 28.7 $97$. $8^{+10.7}_{-10.1}$ $0.86\pm 0.15\pm 0.12$ $1.19\pm 0.61\pm 0.12$ $0.42\pm 0.21\pm 0.20$ 1.8
$3685.26\pm 0.20$ 1.336 26.0 $256$. $2^{+16.8}_{-16.2}$ $0.76\pm 0.10\pm 0.12$ $0.16\pm 0.32\pm 0.12$ $0.57\pm 0.15\pm 0.16$ 0.3
$3686.50\pm 0.20$ 1.338 25.1 $283$. $0^{+17.8}_{-16.8}$ $0.66\pm 0.12\pm 0.12$ $0.02\pm 0.27\pm 0.12$ $0.70\pm 0.15\pm 0.15$ 0.0
$3691.36\pm 0.20$ 1.344 69.4 $77$. $5^{+9.8}_{-8.8}$ $0.16\pm 0.23\pm 0.12$ $1.29\pm 0.54\pm 0.12$ $1.31\pm 0.30\pm 0.16$ 2.8
$3709.76\pm 0.20$ 1.360 70.3 $22$. $0^{+5.8}_{-4.7}$ $0.01\pm 0.40\pm 0.12$ $-2.64\pm 0.60\pm 0.12$ $1.55\pm 0.62\pm 0.19$ 2.0
AVG. -- -- -- $0.69\pm 0.05\pm 0.05$ $0.14\pm 0.16\pm 0.05$ $0.72\pm 0.08\pm 0.07$-
The comparison between this work and previous measurements BESIII:2023sgt ; BESIII:2023ynq is also provided in figure 4. Here, $\Delta\Phi$ is less than zero for $J/\psi$ decay and $\sqrt{s}=3.71$ GeV, and greater than zero at other energy points, which implies that there may be at least one $\Delta\Phi=0$ rad between these energy points. Such an evolution is important input for understanding its asymptotic behavior and the dynamics of baryons similar to the $\Lambda$ hyperon Mangoni:2021qmd .
@@
$M(\cos\theta_{\Sigma^{+}})=\frac{\alpha_{\Sigma^{+}}-\alpha_{\bar{\Sigma}^{-}}}{2}\frac{1+\eta\cos^{2}\theta_{\Sigma^{+}}}{3+\eta}P_{y}(\theta_{\Sigma^{+}}).$ (8)